OpenStudy (clara1223):

f(x)=(x^2)+1+(1/(x^4)), c=1; (a) find the slope of the line that is tangent to the graph of f(x) at the given point x=c; (b) write an equation of the tangent line to the graph of f(x) at the point (c,f(c))

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[f(x)=x^2+1+\frac{1}{x^4}\]?

OpenStudy (clara1223):

yep!

OpenStudy (misty1212):

do you know the derivative?

OpenStudy (misty1212):

lol another county heard from you got \(f'(x)\) if not let me know

OpenStudy (clara1223):

I get that the derivative is 2x-(4/(x^5))

OpenStudy (clara1223):

yeah this is calc

OpenStudy (misty1212):

ok looks good plug in 1 what do you get?

OpenStudy (clara1223):

-2

OpenStudy (misty1212):

ok so \(m=-2\) next i guess find \(f(1)\)

OpenStudy (misty1212):

oh, the answer to question 1 was \(-2\) now we can find the equation of the tangent line, once we have the point

OpenStudy (clara1223):

how do you construct the equation of the tangent line? I know in class we use y-y1=m(x-x1) and that m=f'(c) but what are y1 and x1?

OpenStudy (anonymous):

\[x_1=1,y_1=f(1)\]

OpenStudy (clara1223):

Oh! duh! (c,f(c)). the question told me that. thanks!!